OpenStudy (anonymous):

Tell me if I did this right 9(w-3)=36 w=7

OpenStudy (vijay):

right

OpenStudy (anonymous):

9w - 27 =36 9w = 63 w = 7

OpenStudy (anonymous):

w-3 =36/9 w-3=4 w=4+3 w=7